Show that gauss' law can be used to have an equation that shows the capacitance of two parallel playes in a vacuum.


1
Expert's answer
2021-02-07T19:09:53-0500
"\int EdA=\frac{Q}{\epsilon_0}\to E=\frac{Q}{A\epsilon_0}"

"V=Ed\\C=\frac{Q}{V}=\frac{Q}{Ed}=\frac{QA\epsilon _0}{Qd}
\\C=\frac{A\epsilon _0}{d}"